Target Corporation Income Statement

Wal-Mart Stores,Inc. Date of year

Net sales

61,471

374,526

Cost of goods sold

41,895

286,515

Selling and administration expenses

16,200

70,847

Interest expense

647

1,798

Other income (expense)

1,896

4,273

Income tax expense

1,776

6,908

Net Income

2,849

12,731

Balance sheet data (end of year)

Current assets

18,906

47,585

noncurrent assets

25,654

115,929

total assets

44,560

163,514

current liabilities

11,782

58,454

long-term debt

17,471

40,452

total stockholder's equity

15,307

64,608

total liabilites and stockholders equity

44,560

163,514

Beginning-of-year balances

total assets

37,349

151,587

total stockholder's equity

15,633

61,573

current liabilities

11,117

52,148

total liabilities

21,716

90,014

Other Data

Average net receivables

7,124

3,247

average inventory

6,517

34,433

net cash provided by operating activities

4,125

20,354

Please complete this Project 5 by correctly entering and formatting the selected financial data of Target and Walmart for a recent year as presented (in millions) in the first sheet of this Excel spreadsheet in 3 columns in this current sheet of this Excel spreadsheet provided. I have partially completed the Excel spreadsheet. You need to fill out the Walmart financial data in the Walmart column D, which already has been formatted like the Target column B. Use formulas to calculate subtotals for Gross profit, Operating income, etc. like the Target column I. Please do not add zeros or pennies.

Instructions:

(a)    For each company, compute the 11 ratios using formulas in the columns to the right of the financial data and ratio names. Ratio formulas are in Project 5 Ratios sheet in this Excel spreadsheet. Round all calculated ratios and percentages to 1 decimal place. There is no Preferred Stock so there are no Preferred Stock Dividends. Average collection period is also Days in receivables.

(b)   Compare the overall liquidity, profitability, and solvency of the two companies by providing a specific analysis of these three categories of ratios.

(c)    I have set a Portrait Print Area. Please Print Preview to make sure that the 1 Excel sheet still prints on 1 page.

Target

Walmart

Corporation

Stores, Onc.

Target

Walmart

Income Statement Data for Year

Ratios

Corporation

Stores, Onc.

Net Sales

$           61,471

1

Current ratio

Cost of Goods Sold

               41,895

2

Receivables turnover (Times per year)

Gross Profit

$           19,576

3

Average collection period (Days)

Operating Expenses

4

Inventory turnover (Times per year)

    Selling and administrative expenses

               16,200

5

Days in Inventory (Days)

Opereating Income

$              3,376

6

Profit margin (%)

Non-operating items

7

Asset turnover

    Interest expense

$                 647

8

Return on assets (%)

    Other income (expense)

                 1,896

9

Return on common stockholders' equity (%)

Total non-operating items

$              1,249

10

Debt to total assets (%)

Income before income tax

$              4,625

11

Times interest earned

Income tax expense

                 1,776

Net Income

$              2,849

Balance Sheet Data (End of Year)

Current asssts

$           18,906

Noncurrent assets

               25,654

Total assets

$           44,560

Current liabilities

$           11,782

Long-term debt

               17,471

Total liabilities

$           29,253

Total stockholders' equity

$           15,307

Total liabilities and stockholders' equity

$           44,560

Beginning-of-Year Balances

Total assets

$           37,349

Total stockholders' equity

               15,633

Current liabilities

               11,117

Total liabilities

               21,716

Other Data

Average net receivables

$              7,124

Average Inventory

                 6,517

Net cash provided by operating activities

                 4,125

Liquidity:

Answer:

Profitability:

Answer:

Solvency:

Answer:
Answer #1

Income Statement

Target corp($) Walmart($)
Net sales 61471 374526
Cost of goods sold 41895 286515
Gross Profit 19576 88011
Operating Exp
Selling and administration exp 16200 70847
Operating income 3376 17164
Non operating exp
Intt exp 647 1798
Other income 1896 4273
Total non operating items 1249 2475
Income before taxes 4625 19639
Tax exp 1776 6908
Net income 2849 12731
Target Corp($) Walmart ($)
Current assets 18906 47585
Non current assets 25654 115929
Total assets 44560 163514
Current liab 11782 58454
Long term debt 17471 40452
Total liab 29253 98906
Equity 15307 64608
Total equity and liab 44560 163514
Beginning values
Total assets 37349 151587
Total stockholders equity 15633 61573
Current liab 11117 52148
Total liab 21716 90014
Other data
Average Receivable 7124 3247
Average inventory 6517 34433
Net cash from operating operations 4125 20354
Ratios formula target corp Walmart
1.current ratio current assets/current liab 18906/11782=1.6 47585/58454=0.81
2. Receivable turnover ratio Sales/Average Receivable 61471/7124=8.63 times

374526/3247=115.34 times

3. Average collection period 365/Receivable turnover ratio 42.29 days 3.16 days
4.Inventory turnover sales/ average inventory 61471/6517=9.43 times 374526/34433=10.88 times
5. Days in Inventory 365/turnover ratio 38.71 days 33.55 days
6. Profit margin net income/ sales*100 2849/61471*100= 4.63% 12731/37452*100=33.99%
7.assets turnover ratio sales/ beg assets 61471/37349=1.64 times 374526/151587=2.47 times
8.return on assets net income/assets*100 2849/44560*100=6.39% 12731/163514*100=7.78%
9. Return on stockholders equity net income/ stockholders equity 2849/15307*100= 18.61% 12731/163514*100=7.78%
10. Debt to total assets Long term debt/total assets*100 17471/44560*100=39.21% 40452/163514*100=24.74%

11.time interest income before Intt & taxes/Int

Earned=

Target corp= 3376/647=5.22 times

Walmart= 17164/1798=9.55 times

Please give your feedback!! Happy Learning :)